Sun radiates energy in all direction the average energy received at earth is 1.4 KW/m² . Average distance betn earth and sun is 1.5×10¹¹ m. If the energy is released by conservation of mass and energy then mass lost by sun per day is approximately is ............................kg.

Grade:12th pass

1 Answers

Vikas TU
7 years ago
 A = 4•π•r^2 = 4•π•(1.5•10^11)^2 = =9•π•10^22 m^2. 

b. Power density is p =1400 W/m^2. 
The total power is 
P = p•A =1400 • 9•π•10^22 = 
=3.96•10^26 W. 

c. Stefan-Boltzmann Law 
R = σ•T^4, 
R =P/A1 = P/(4• π•r1^2), is the black body irradiance, 
σ =5.67• 10^-8 W•m^-2•K^-4 is the Stefan-Boltzmann constant, 
T =[P/(A1• σ)]^1/4 ={3.96•10^26/[4• π•(6.96• 10^8)^2]}^1/4 = 5800 K.